u - 1.96t = 10
u + 1.96t = 30

Answer by onlinetutor365.com(14)

u - 1.96t = 10 --(1)
u + 1.96t = 30 --(2)
Add (1)+ (2) ==> 2u=40
u=40/2
u=20
Put u value into (1) ==> 20-1.96t=10
-1.96t=10-20
-1.96t=-10
1.96t=10
t=10/1.96
t=5.102
Therefore, u=20 and t=5.102